OpenStudy (ibbutibbu.):

Solve for x: -3|x + 7| = -12

OpenStudy (ibbutibbu.):

one more

OpenStudy (ibbutibbu.):

And just one more

OpenStudy (neer2890):

The value of mod(x+7) depends upon the value of x. When the value of x is less then -7, |x+7| becomes -(x+7) otherwise , it's (x+7) itself. we can define it as a function: \[f(x)= \left(\begin{matrix} -(x+7),x<-7\\ (x+7),x \ge -7\end{matrix}\right)\] so we can solve it as: \[-3|x+7|=-12\] \[|x+7|=3\] Case1: when -(x+7)=3,x<-7 -x-7=3 x=-10 Case 2: when x+7=3,x>=-7 x+7=3 x=-4

OpenStudy (ibbutibbu.):

but... these are the answer choices- x = -3, x = -11 x = , x = - x = -3, x = 11 No solution

OpenStudy (neer2890):

what;s 2nd option?

OpenStudy (ibbutibbu.):

oh sry it is x=5/3 x= -19/3

OpenStudy (neer2890):

sorry, i've got a mistake -3|x+7|=-12 |x+7|=4 instead of |x+7|=3 Case1: -(x+7)=4 -x-7=4 x=-11 Case2: (x+7)=4 x=-3 so, the answer is x=-3 ,x=-11

OpenStudy (ibbutibbu.):

ok thank you

OpenStudy (neer2890):

you're welcome..:) try similar approach for attachment 1.